How much heat is required to raise the temperature of 85 grams of water from 280 K to 342 K?
  1. 5270 J
  2. 355 J
  3. 355 J
  4. 151 J
  5. 22029 J

Grade:11

3 Answers

Shivangi Khatter
5 years ago
dear student
the correct answer is(A)
Heat= mc deltaT = 85 1 62= 5270 J
Arun
5 years ago
Dear student
 
as 
Q = ms (delta T) =85 *10^-3 * 4.18*10^3 * 62 = 22029 J
 
Hope it helps
 
Regards
Arun
APARNA SHAJI
4 years ago
THE COORECT ANSWER IS A AS 
HEAT= MC DELTA T AND C FOR WATER IS 1.
SO 85*1*62=5270 J
PLS APPROVE IF IT HELPED U